### v1

@@ Abstract
Large Language Models (LLMs) have made notable progress in mathematical reasoning, yet they often rely on single-paradigm reasoning that limits their effectiveness across diverse tasks. In this paper, we introduce Chain-of-Reasoning (CoR), a novel unified framework that integrates multiple reasoning paradigms—Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR)—to enable synergistic collaboration. CoR generates multiple potential answers using different reasoning paradigms and synthesizes them into a coherent final solution. We propose a Progressive Paradigm Training (PPT) strategy that allows models to progressively master these paradigms, culminating in the development of CoR-Math-7B. Experimental results demonstrate that CoR-Math-7B significantly outperforms current SOTA models, achieving up to a $41.0\%$ absolute improvement over GPT-4 in theorem proving tasks and a $7.9\%$ improvement over RL-based methods in arithmetic tasks. These results showcase the enhanced mathematical comprehensive ability of our model, achieving significant performance gains on specific tasks and enabling zero-shot generalization across tasks.
Figure 1: A comprehensive comparative analysis of CoR-Math-7B and baseline models across challenging mathematical tasks. (a) demonstrates the effectiveness of CoR-Math-7B (zero-shot) in theorem proving tasks, specifically emphasizing its advantage over baseline and their fine-tuned variants over SFT and RL. (b) shows a resource-efficiency analysis for arithmetic computation tasks, plotting model performance vs. SFT data size, in which CoR-Math-7B attains the optimal resource efficiency, achieving near-optimal zero-shot performance.
@@ 1 Introduction
While LLMs have demonstrated strong performance in solving mathematical tasks (Feigenbaum et al., 1963; Hosseini et al., 2014), advanced open-source reasoners still struggle with solving comprehensive mathematical problems, including both arithmetic computation and theorem proving.
@@
Existing works (Xin et al., 2024; Yang et al., 2024; Wu et al., 2024) are often trained on specific tasks, aiming to enhance their ability to independently derive answer based on specific structured knowledge representation. This representation is known as the reasoning paradigm, involving Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR), as illustrated in Figure 2 (a). Specifically, NLR leverages natural language text for reasoning based on human common sense and semantic context, with explicit step-by-step natural language explanations (Wei et al., 2022). AR leverages code to focus on the computer’s operations and execution process, performing inference on the final target, such as the generation of Python code for compilation (Chen et al., 2023; Gao et al., 2023) to obtain the final answer. SR utilizes logical symbols and axiomatic systems as the fundamental tools for rigorously formalized reasoning, with current methods (Xin et al., 2024; Huang et al., 2024; Wu et al., 2024) considering numerous symbolic trajectories through a tree-based search process for theorem proving. However, these methods focus on improving single-paradigm reasoning and overlook the potential of collaboration among multiple paradigms, which restricts their single-task performance and hinders their cross-task generalization.
@@
Researchers have explored various strategies to tackle these challenges. To improve the single-task performance, some works have explored integrating tools to overcome the limitations of single-paradigm reasoning (Gou et al., 2024; LI et al., 2024), as shown in Figure 2 (b). These methods combine natural language with code-based algorithms (tools) to enable interleaved reasoning and generate a final answer. Although these methods acknowledge the potential benefits of integrating different reasoning paradigms, they often rely on a single paradigm to complete the reasoning process. They overlook the possibility that the second paradigm could independently complete the reasoning, thereby constraining its overall potential. Besides, in an effort to overcome the challenge of achieving cross-task generalization, several studies (Shao et al., 2024; Huang et al., 2024) incorporate data samples from diverse task types into large-scale training datasets, such as those drawn from theorem proving tasks that focus exclusively on SR solutions, or from arithmetic problems that emphasize NLR solutions. Although models trained on such data are capable of cross-task reasoning, they still rely on demonstrations for effective transfer.
@@
To address these limitations, we propose a novel unified reasoning framework, Chain-of-Reasoning (CoR), which chains NLR, AR, and SR together to generate synergistic benefits. As illustrated in Figure 2 (c), CoR enables multi-paradigm reasoning for a given problem, which applies different reasoning paradigms to derive multiple potential answers and are then summarized into a final answer. The framework allows the model to continue reasoning using additional paradigms based on previously generated ones, facilitating collaboration among paradigms to enhance individual task performance. Moreover, CoR implements unified multi-paradigm reasoning across different tasks to obtain the required answers, thereby achieving zero-shot reasoning across tasks. In detail, by adjusting the prompts, the depth of reasoning can be varied, which improves the model’s adaptability to diverse tasks. By constructing such multi-paradigm reasoning paths, we create a training dataset named MPM, comprising $167$k reasoning paths. Furthermore, we propose a novel Progressive Paradigm Training (PPT) approach, which enables models to progressively master multiple reasoning paradigms, resulting in CoR-Math-7B.
@@
We evaluated CoR-Math-7B on three representative and challenging mathematical reasoning benchmarks that effectively demonstrate comprehensive reasoning capabilities, encompassing both arithmetic computation and theorem proving. These benchmarks include MATH (Hendrycks et al., 2021), GSM8k (Cobbe et al., 2021), and MiniF2F (Zheng et al., 2022). Our results show that LLMs equipped with CoR significantly outperform current SOTA baselines. In the theorem proving task as shown in Figure 1 (a), CoR-Math-7B surpasses GPT-4o’s (Hurst et al., 2024) zero-shot performance by a remarkable absolute $41.0\%$ and outperforms all other few-shot reasoners. In arithmetic tasks, our CoR-Math-7B first performs NLR and SR as a rehearsal, which enhances the accuracy of AR and demonstrates the effectiveness of synergy. For example, CoR-Math-7B achieves an absolute improvement of $7.9\%$ over Reinforcement Learning (RL) methods (Shao et al., 2024) that utilize the same base model on the MATH dataset. Furthermore, as shown in Figure 1 (b), CoR-Math-7B efficiently utilizes resources to surpass the optimal performance curve of single-paradigm approaches. Compared to mainstream methods that perform extensive searches within a single paradigm, our CoR framework achieves better results by extending test-time inference across multiple paradigm spaces. These results show that our method achieves comprehensive mathematical problem solving abilities though multi-paradigm reasoning, requiring fewer training data and reducing reasoning costs.
@@ 2 Related Work
Reasoning Paradigms in LLMs. Recent advancements in LLMs have primarily concentrated on enhancing single-paradigm reasoning, where each paradigm serves as a distinct medium for knowledge representation and logical inference. NLR utilizes human-readable language to solve commonsense reasoning and step-by-step deductions (Wei et al., 2022; Yao et al., 2023; Zhou et al., 2023; Besta et al., 2024; Sel et al., 2024). For precise numerical computations, AR is to generate executable code, such as Python scripts, to perform accurate calculations through execution (Chen et al., 2023; Rozière et al., 2023). In the realm of theorem proving, SR leverages logical symbols and axiomatic systems to ensure rigorous and verifiable proofs (Xin et al., 2024). Furthermore, some approaches leverage external tools within the single-paradigm to enhance reasoning capabilities (Gou et al., 2024), such as using calculators or code interpreters, while maintaining the core reasoning process within a single paradigm. While these paradigms excel within their specific areas, they often struggle with cross-domain generalization and dynamic environments. To address these issues, the CoR framework facilitates collaboration between various reasoning paradigms to enable zero-shot multitasking generalization.
Mathematical Problem Solving with LLMs. Advanced research demonstrates the significant potential of LLMs in solving mathematical problems (Zhu et al., 2023). Several studies have developed unified mathematical solvers by synthesizing mathematical data, capable of addressing various challenges such as arithmetic calculations and theorem proving (Huang et al., 2024; Shao et al., 2024). However, recent research tends to focus on optimizing single paradigms. For instance, in the field of arithmetic computation, there is a tendency to incorporate programming tools to aid natural language mathematical reasoning (Gou et al., 2024). In theorem proving, researchers not only utilize specialized annotated data, including pre-training data and reinforcement learning reward data, but also employ tree-based search methods to generate numerous possible solutions (Ying et al., 2024b; Xin et al., 2024). These methods either fail to fully consider the importance of complete reasoning or rely heavily on large-scale search within the solution space to fully explore the performance improvement potential of a single paradigm. To address these issues, we introduce the MPM dataset and present the CoR-Math-7B model, trained on it, which leverages the synergistic effects of multi-paradigm reasoning to navigate towards the optimal solution within the expanded multi-paradigm solution space.
@@ 3.1 Overview
Our CoR framework aims to enable LLMs to perform a series of multi-paradigm reasoning on any type of mathematical problem, ultimately arriving at a solution. Specifically, given a mathematical problem $x$, LLMs ($\mathbb{P}$) can infer the result $y$ by following multiple reasoning paradigms, where each reasoning paradigm $\tau$ includes $n$ reasoning paths $\{rp_{1},...,rp_{n}\}$. To formally represent this single-paradigm scenario, we can formulate it as $y\sim\mathbb{P}(y|x,\tau)$. To simplify the process for each reasoning paradigm, we set $n=1$ as default (Details in Section 6).
Inspired by recent works (Wei et al., 2022) in encouraging step-by-step reasoning, we introduce CoR, which extends from a single paradigm to three paradigms $\Gamma=(\tau_{1},\tau_{2},\tau_{3})$. For generating multiple chained reasoning paradigms for a given problem, CoR follows the steps outlined below. The reasoning process begins with the initialization of the problem $x$ and the first reasoning paradigm $\tau_{1}$. Subsequently, each paradigm $\tau_{i}$ in the sequence $\Gamma$ is generated based on the problem $x$ and the previously established paradigms $\tau_{1},...,\tau_{i-1}$, as represented by $\tau_{i}\sim\mathbb{P}(\tau_{i}|x,\tau_{1},...,\tau_{i-1})$. Finally, the outcomes from all reasoning paradigms are aggregated to derive the final result $y$, expressed as $y\sim\mathbb{P}(y|x,\tau_{\text{NLR}},\tau_{\text{SR}},\tau_{\text{AR}})$, considering three paradigms: NLR, SR, and AR. In detail, $\tau_{\text{SR}}$ and $\tau_{\text{AR}}$ include both the complete reasoning paths and the interaction processes with tool outputs. In our experiments, $\tau_{\text{SR}}$ utilizes the Lean prover, and $\tau_{\text{AR}}$ leverages a Python compiler as the algorithmic interactive tool. CoR possesses the capacity for broad application in complex reasoning scenarios. In the context of mathematical reasoning, it manifests as CoR-Math, with potential for future adaptations such as CoR-Code.
According to this framework, as shown in Figure 3, the training pipeline consists of three key phases: (a) collecting a Multi-Paradigm Mathematical (MPM) dataset, with deep reasoning paths based on multiple reasoning paradigms; (b) introducing a Progressive Paradigm Training (PPT) strategy to shape the model’s reasoning behavior, gradually mastering more reasoning paradigms; and (c) leveraging the trained LLM for zero-shot inference, allows multi-paradigm reasoning with adaptable depth, and uses Sequential Multi-Paradigm Sampling (SMPS) to explore a diverse solution space.
@@ 3.2 Collecting Dataset
[…]
Stage 1: Reconstructing and Extending. We introduce a universal text template designed for multi-paradigm reasoning (Details in Section A.1). This template standardizes the positioning of distinct reasoning paradigms and defines the relationships between them. It accommodates various reasoning depths and supports flexible combinations of different reasoning paradigms. As shown in Figure 2 (c), Problem 1 demonstrates an instance incorporating both NLR and SR paradigms, while Problem 2 showcases an instance that integrates three reasoning paradigms. To ensure data integrity and prevent potential biases, we pre-process the Numina-TIR (LI et al., 2024) and Lean-Workbook (Ying et al., 2024a) datasets through a two-step process. First, samples without corresponding solutions are removed. Second, we further reconstruct and extend these datasets by leveraging powerful LLMs $\mathcal{G}$, such as GPT-4o (Hurst et al., 2024). These models generate missing reasoning paradigms $\tau_{g}$ and refine existing ones $\tau^{\prime}$, ensuring comprehensive coverage and logical consistency. To effectively guide the processes of augmentation and refinement, we develop tailored prompts $p_{s}$ for each seed dataset. Examples of these prompts are detailed in Section A.2. The procedure can be described in:
@@
where $\oplus$ means concatenation. After that, we conduct a manual review of all samples. Given that alternative approaches are readily verified through external tools, we focus on the accuracy of the NLR and AR. This approach considerably lowers the requisite skill level of annotators. Furthermore, to avoid data leakage, we calculate the Levenshtein distance (Miller et al., 2009) between problems in the training and test samples, while details are provided in the Section B.1. As a result, this phase yields the MPM-raw dataset, comprising approximately $285,000$ synthetic samples.
Stage 2: Revising. The MPM-raw dataset interacts with the Lean prover to verify proof steps, guiding the filtering and modification of reasoning paths. The process submits the proof $\tau_{\text{SR}}$ to the prover. Once the prover successfully completes the proof without returning any errors, the entire multi-paradigm reasoning path is collected to the MPM dataset. Otherwise, the error information $\varepsilon$ returned by the prover is fed into a revising model $\mathbb{P}_{R}$. Furthermore, this model generates a revised proof $\tilde{\tau}_{\text{SR}}$ based on a prompt $p_{\varepsilon}$. The relationship can be expressed as:
@@
where the revised proof $\tilde{\tau}_{\text{SR}}$ is then resubmitted to the Lean prover for verification. This iterative process continues for a maximum of $64$ iterations or until the prover verifies $\tilde{\tau}_{\text{SR}}$ as correct. In detail, we utilize DeepSeek-Prover-V1.5 (Xin et al., 2024) as the revising model.
As a result, the MPM dataset comprises $82,770$ problems and $167,412$ multi-paradigm reasoning solutions.

### v4

@@ Abstract
Large Language Models (LLMs) have made notable progress in mathematical reasoning, yet they often rely on single-paradigm reasoning that limits their effectiveness across diverse tasks. In this paper, we introduce Chain-of-Reasoning (CoR), a novel unified framework that integrates multiple reasoning paradigms — Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR) — to enable synergistic collaboration. CoR generates multiple potential answers using different reasoning paradigms and synthesizes them into a coherent final solution. We propose a Progressive Paradigm Training (PPT) strategy that allows models to progressively master these paradigms, culminating in the development of CoR-Math-7B. Experimental results demonstrate that CoR-Math-7B significantly outperforms current SOTA models, achieving up to a $41.0\%$ absolute improvement over GPT-4o in theorem proving tasks and a $15\%$ improvement over RL-based methods on the MATH benchmark in arithmetic tasks. These results show the enhanced mathematical comprehensive ability of our model, enabling zero-shot generalization across tasks. The code is available at https://github.com/microsoft/CoR.
[…]
[Uncaptioned image] Chain-of-Reasoning: Towards Unified Mathematical Reasoning in Large Language Models via a Multi-Paradigm Perspective
[…]
@@ 1 Introduction
Figure 1: A comprehensive comparative analysis of CoR-Math-7B and baseline models across mathematical tasks. (a) shows the effectiveness of CoR-Math-7B (zero-shot) in theorem proving tasks. (b) shows a resource-efficiency analysis for arithmetic computation tasks, where CoR-Math-7B achieves optimal resource efficiency and near-optimal zero-shot performance.
While LLMs have shown strong performance in solving mathematical tasks (Feigenbaum et al., 1963; Hosseini et al., 2014), advanced open-source reasoners still struggle with solving comprehensive mathematical problems, including both arithmetic computation and theorem proving.
Existing works (Xin et al., 2024; Yang et al., 2024; Wu et al., 2024; Zhang et al., 2024) are often trained on specific tasks, aiming to enhance their ability to independently derive answers based on specific structured knowledge representation. This representation is known as the reasoning paradigm, involving Natural Language Reasoning (NLR), Algorithmic Reasoning (AR), and Symbolic Reasoning (SR), as depicted in Fig.˜2 (a). Specifically, NLR uses natural language for reasoning via common sense and semantic context, providing explicit step-by-step explanations (Wei et al., 2022). AR leverages code to emphasize computational operations and execution processes, such as generating Python code for execution (Chen et al., 2023; Gao et al., 2023). SR uses logical symbols and axiomatic systems for formalized reasoning, with recent methods (Xin et al., 2024; Huang et al., 2024; Wu et al., 2024) considering numerous symbolic trajectories via tree-based search for theorem proving. However, these methods mainly focus on optimizing single-paradigm reasoning, creating models that demonstrate asymmetrical performance across different mathematical tasks. For instance, models specialized in NLR may exhibit deficiencies in theorem proving and vice versa. This fragmented paradigm constrains the upper-bound performance on individual tasks and undermines the model’s capacity for cross-paradigm generalization.
@@
Researchers have explored various strategies to tackle these challenges. To improve the single-task performance, some works integrate tools to overcome the limitations of single-paradigm reasoning (Gou et al., 2024; LI et al., 2024), as shown in Fig.˜2 (b). While recognizing the benefits of combining reasoning paradigms, they still rely on a single paradigm, neglecting the possibility that the second paradigm could independently complete the reasoning, thereby constraining overall potential. Besides, to improve cross-task generalization, several studies (Shao et al., 2024; Huang et al., 2024) incorporate diverse task samples into large-scale datasets, such as those drawn from theorem proving tasks that focus exclusively on SR solutions, or from arithmetic problems that emphasize NLR solutions. Although models trained on such data are capable of cross-task reasoning, they still rely on demonstrations for effective transfer.
@@
To address these limitations, we propose Chain-of-Reasoning (CoR), a unified framework integrating NLR, AR, and SR to produce synergistic benefits. As shown in Fig.˜2 (c), CoR enables multi-paradigm reasoning by applying different reasoning paradigms to derive multiple potential answers, which are then summarized into a final solution. This framework allows iterative reasoning across paradigms, leveraging prior results to enhance single-task performance. Moreover, CoR unifies multi-paradigm reasoning across tasks, enabling zero-shot generalization. Adjusting prompts to control reasoning depth improves adaptability to diverse tasks. As a result, we introduce Multi-Paradigm Mathematical (MPM), a dataset comprising 167167k reasoning paths, and propose Progressive Paradigm Training (PPT), a method enabling models to progressively master multiple reasoning paradigms, leading to CoR-Math-7B.
@@
We evaluate CoR-Math-7B on five challenging mathematical reasoning benchmarks, covering both arithmetic computation and theorem proving. Our results show that LLMs equipped with CoR significantly surpass current SOTA baselines. In theorem proving (Fig.˜1 (a)), CoR-Math-7B improves zero-shot performance over GPT-4o (Hurst et al., 2024) by $41.0\%$, outperforming all few-shot reasoners. For arithmetic tasks, CoR-Math-7B achieves a $24.2\%$ absolute improvement over GPT-4 on MATH. Furthermore, as shown in Fig.˜1 (b), CoR-Math-7B efficiently utilizes resources to surpass the optimal performance curve of single-paradigm approaches. Unlike mainstream methods that conduct extensive searches within a single paradigm, CoR enhances test-time inference in multiple paradigms. These results show that CoR can solve comprehensive mathematical problems through multi-paradigm reasoning, requiring less training data and lowering reasoning costs.
@@ 2 Related Work
Reasoning Paradigms in LLMs. Recent advancements in LLMs focus on single-paradigm reasoning, with each paradigm representing a distinct method for knowledge representation and logical inference. NLR uses human-readable language for commonsense reasoning and step-by-step deductions (Wei et al., 2022; Yao et al., 2023; Zhou et al., 2023; Besta et al., 2024; Sel et al., 2024), AR generates executable code for precise calculations (Chen et al., 2023; Rozière et al., 2023), and SR utilizes logical symbols for theorem proving (Xin et al., 2024). Some methods enhance reasoning by integrating external tools like calculators or code interpreters, yet they remain within a single paradigm (Gou et al., 2024). While effective in specific tasks, they struggle with cross-domain generalization and dynamic environments. To address these limitations, the CoR framework enables collaboration across reasoning paradigms, enabling zero-shot multitask generalization.
Mathematical Problem Solving with LLMs. Advanced research highlights the potential of LLMs in solving mathematical problems (Zhu et al., 2023; Lin et al., 2024b; Luo et al., 2025). Several studies have developed unified solvers by synthesizing mathematical data to address challenges like arithmetic computation and theorem proving (Huang et al., 2024; Shao et al., 2024). However, most approaches focus on optimizing a single paradigm. For instance, in arithmetic computation, tools are integrated to assist natural language reasoning. (Gou et al., 2024). In theorem proving, models rely on specialized data like pre-training data and Reinforcement Learning (RL) reward data, and tree-based search methods to generate numerous possible solutions (Ying et al., 2024b; Xin et al., 2024). These approaches either neglect complete reasoning or depend on large-scale search within the solution space, limiting performance gains within a single paradigm. To address these, we introduce the CoR-Math-7B model with complete reasoning processes to explore an expanded multi-paradigm solution space.
@@ 3.1 Overview
CoR aims to enable LLMs to perform a series of multi-paradigm reasoning on any type of mathematical problem, ultimately arriving at a solution. Specifically, given a mathematical problem $x$, LLMs ($\mathbb{P}$) can infer the result $y$ by following multiple reasoning paradigms, where each reasoning paradigm $\tau$ includes $n$ reasoning paths $\{rp_{1},...,rp_{n}\}$. We represent this single-paradigm scenario as $y\sim\mathbb{P}(y|x,\tau)$. To simplify the process for each reasoning paradigm, we set $n=1$ as default (Details in Appendix˜A).
Inspired by recent works (Wei et al., 2022) in encouraging step-by-step reasoning, we introduce CoR, which extends from a single paradigm to three paradigms $\Gamma=(\tau_{1},\tau_{2},\tau_{3})$. For generating multiple chained reasoning paradigms for a given problem, CoR follows the steps outlined below. The reasoning process begins with the problem $x$ and the first reasoning paradigm $\tau_{1}$. Subsequently, each paradigm $\tau_{i}$ in the sequence $\Gamma$ is generated based on the problem $x$ and the previously established paradigms $\tau_{1},...,\tau_{i-1}$, as represented by $\tau_{i}\sim\mathbb{P}(\tau_{i}|x,\tau_{1},...,\tau_{i-1})$. Finally, the outcomes of all the reasoning paradigms are aggregated to derive the final result $y$, expressed as $y\sim\mathbb{P}(y|x,\tau_{\text{NLR}},\tau_{\text{SR}},\tau_{\text{AR}})$, considering three paradigms: NLR, SR, and AR. In detail, $\tau_{\text{SR}}$ and $\tau_{\text{AR}}$ include both the complete reasoning paths and the interaction processes with tool outputs. In this study, $\tau_{\text{SR}}$ uses the Lean prover, and $\tau_{\text{AR}}$ applies a Python compiler to obtain reasoning results.
In our workflow (Fig.˜3), the training pipeline includes: (a) collecting the MPM dataset with deep reasoning paths; (b) using progressive paradigm training to enhance reasoning across paradigms; and (c) applying the trained LLMs for zero-shot inference with sequential multi-paradigm sampling to explore diverse solutions.
@@ 3.2 Collecting Dataset
To train CoR model, we extend the single-paradigm datasets to incorporate multiple reasoning paradigms, denoted as $\texttt{x},\texttt{NLR},\texttt{SR},\texttt{AR},\texttt{y}>$. As presented in Fig.˜3 (a), the training data collection process involves two stages: (a.1) Reconstructing and Extending, and (a.2) Revising. In the first stage, we reconstruct high-quality open-source mathematical data as seed samples and synthesize additional reasoning paradigms to form the MPM-raw dataset. In the second stage, a theorem prover examines MPM-raw samples, and a mathematical reasoner revises those that fail, compiling all completed data into the final MPM dataset.
Stage 1: Reconstructing and Extending. We introduce a universal text template for multi-paradigm reasoning (details in Sec.˜B.1), which standardizes the placement and relationships of reasoning paradigms. It supports various reasoning depths and flexible combinations of different reasoning paradigms. As shown in Fig.˜2 (c), Problem 1 shows an instance incorporating both NLR and SR paradigms, and Problem 2 integrates three reasoning paradigms. To ensure data integrity and prevent potential biases, we pre-process the Numina-TIR (LI et al., 2024) and Lean-Workbook (Ying et al., 2024a) datasets in two steps. First, samples without corresponding solutions are removed. Second, we further reconstruct and extend these datasets by leveraging powerful LLMs $\mathcal{G}$, such as GPT-4o (Hurst et al., 2024). These models generate missing reasoning paradigms $\tau_{g}$ and refine existing ones $\tau^{\prime}$, ensuring comprehensive coverage and logical consistency. To effectively guide the processes of augmentation and refinement, we develop tailored prompts $p_{s}$ for each seed dataset (examples in Sec.˜B.2). The procedure can be described in:
@@
where $\oplus$ means concatenation. After that, we conduct a manual review of all samples. Given that alternative approaches are readily verified through external tools, we focus on the accuracy of the NLR and AR. This approach considerably lowers the requisite skill level of annotators. To prevent data leakage, we compute the Levenshtein distance (Miller et al., 2009) between training and test problems, as detailed in Sec.˜C.1. As a result, this phase yields the MPM-raw dataset, comprising approximately $285,000$ synthetic samples.
Stage 2: Revising. The MPM-raw dataset interacts with the Lean prover to verify proof steps, guiding the filtering and modification of reasoning paths. The proof $\tau_{\text{SR}}$ is submitted to the prover, and if verification is successful without errors, the entire multi-paradigm reasoning path is collected to the MPM dataset. Otherwise, the error information $\varepsilon$ returned by the prover is fed into a revising model $\mathbb{P}_{R}$. Furthermore, this model generates a revised proof $\tilde{\tau}_{\text{SR}}$ based on a prompt $p_{\varepsilon}$. The relationship can be expressed as:
@@
where the revised proof $\tilde{\tau}_{\text{SR}}$ is then resubmitted to the Lean prover for verification. This iterative process continues for up to 64 iterations or until $\tilde{\tau}_{\text{SR}}$ is verified as correct. In detail, we employ DeepSeek-Prover-V1.5 (Xin et al., 2024) as the $\mathbb{P}_{R}$.
Consequently, the MPM dataset comprises $82,770$ problems and $167,412$ multi-paradigm reasoning solutions.